# arXiv:2602.22437v2[cs.DC] 27 Feb 2026

### v2

@@ 1 Introduction
Large language models (LLMs) have become a transformative technology in everyday applications. Driven by the scaling law [kaplan2020scaling], LLMs now reach billions of parameters and achieve human-level performance. Training such giant models requires parallelization techniques that distribute the model and optimizer states across thousands of GPUs [jiang2024megascale]. Among these, Deepspeed ZeRO [rajbhandari2020zero] or Fully Sharded Data Parallel (FSDP) [zhao2023pytorch, pytorch2024fsdp2, megatron_fsdp] is one of the most fundamental techniques. FSDP is often the first choice because of its efficient yet flexible data-parallel programming paradigm and decoupling from model architecture. When additional scaling is needed [smith2022using, ma2025veomni], FSDP can be combined with other parallelisms.
@@
However, existing FSDP/ZeRO systems struggle with modern structure-aware training. State-of-the-art models use non-element-wise optimizers such as Shampoo [gupta2018shampoo] and Muon [jordan2024muon], and block-wise quantized training like DeepSeek-V3 [liu2024deepseek], all of which require atomic tensor blocks. The core limitation is that existing FSDP frameworks shard parameters, gradients, and optimizer states either element-wise [rajbhandari2020zero, zhao2023pytorch] or row-wise [pytorch2024fsdp2, megatron_fsdp], producing sharding boundaries that often misalign with the required block sizes. Consequently, model developers must intrusively modify their models or optimizers to match tensor boundaries, or system developers must handle complex boundary checks, padding, and additional communication logic.
@@
Beyond inflexibility, current FSDP systems fall short of our production throughput and memory targets, where we aim to extract every bit of hardware efficiency. GPU Memory is the tighter constraint: in shared clusters, jobs run out of memory or will operate at the memory limit incurring expensive device-side frees, prompting over-provisioning that leaves GPU resources wasted. These demands become even more critical when scaling training to over 10K GPUs and trillions of parameters. Few existing FSDP systems can scale to this level while maintaining efficiency. Deepspeed ZeRO [rajbhandari2020zero] pioneered the FSDP research but suffers from fragmented AllGather operations [deepspeed_AG] and inefficient memory management [fsdp_record_stream]. PyTorch FSDP1 [zhao2023pytorch] addresses some AllGather inefficiency, but incurs slow ReduceScatter [fsdp1_reduce] and does not solve memory overhead [fsdp_record_stream]. PyTorch FSDP2 [pytorch2024fsdp2] improves memory management [per_parameter_shard_rfc] but introduces high tensor copy overhead. Meanwhile, both FSDP1 and FSDP2 suffer from slow collectives due to unaligned communication buffer [wu2025terabyte, nccl16byte]. Megatron-FSDP [megatron_fsdp] further improves performance but requires extra padding, increasing both communication and memory costs.
To this end, we reinvent PyTorch FSDP2 and present veScale-FSDP, combining both flexibility and performance at scale:
@@
$\triangleright$ For performance, veScale-FSDP introduces a planning algorithm that rearranges RaggedShard tensors to maximize communication efficiency while respecting their desired sharding granularities. We formulate planning as a NP-hard optimization problem and use practical polynomial-time heuristics that achieve high-quality solutions in practice.
$\triangleright$ veScale-FSDP further provides a high-performance primitive, Distributed Buffer (DBuffer), that backs RaggedShard tensors with slices of a global buffer, not only enabling zero-copy access and minimal communication overhead but also reducing memory fragmentation via batched memory allocations.
@@
RaggedShard code is open sourced at https://github.com/volcengine/veScale.
@@ 2.1 Structure–Aware Training
Structure–aware training is the core technique behind the top-tier models such as Gemini [team2024gemini] and DeepSeek [liu2024deepseek], and becomes increasingly important, including:
Matrix Optimizers.   Matrix-based optimizers such as Shampoo [gupta2018shampoo] and Muon [jordan2024muon] can deliver faster convergence rate. The calculation is conducted on the matrix with the original 2D shape, requiring the full matrix to be present locally on each device and then be computed only on chosen devices.
Block-wise Quantization.   Training with quantized model [liu2024deepseek] and optimizer states [dettmers20218] is widely used to improve system efficiency. Block-wise quantization is one of the prevailing techniques to preserve both quality and efficiency of training, but requires slicing the tensors into 2D blocks for the calculation of the scaling factors. Sharding parameters without care would end up in sharded blocks across devices, incurring high complexity in either model design or system development.
@@ 2.2 DTensor and JaggedTensor
Figure 1: The DTensor for flexible communication and computation. Here shows an example of DTensors executing a sharded matrix multiplication on a device. The darken part in each DTensor indicates the materialized local tensor on this device.
Distributed Tensor (DTensor)   [torch-dtensor, li2025vescale] is a promising primitive of PyTorch that provides the opportunity towards structure–aware training. It represents a global tensor distributed across devices, where each device holds a local sharded tensor. DTensor supports three sharding formats (placement): Shard(dim) that evenly shards a global tensor along a tensor dimension, Replicate that replicates the global tensor, and Partial. It also enables users to switch between these placement via redistribute with implicit collective communications. Additionally, DTensors can be computed directly by operators like matmul, as shown in Figure 1. However, a fundamental limitation prevents structure-aware training: the Shard format cannot represent the block-wise sharding needed for quantization or the uneven sharding required by matrix optimizers.
JaggedTensor/NestTensor on a single device   [jaggedtensor, nestedtensor, raggedtensors] are PyTorch/TensorFlow primitives that represent single-device tensors whose last dimension is “jagged”. For example, a 2D tensor in which each row may have a different length. While these primitives still fail to represent the block-level granularity as the atomic unit, they offer a useful hint for how veScale-FSDP can support structure–awareness in distributed training setting.
@@ 2.3 ZeRO and FSDPs
DeepSpeed ZeRO   [rajbhandari2020zero] pioneered this line of FSDP research. Its core idea is to concatenate a layer of tensors (parameters, gradients, and optimizer states) and then shard each concatenated tensors across devices, where tensors can be irregularly sharded across device boundary. ZeRO only unshards a layer using AllGather before forward and backward pass, and reduces the layer gradients using ReduceScatter back to different devices. Such sharding design is limited in element-wise plain tensor and cannot support structure-aware training.
FullyShardedDataParallel (FSDP1)   [zhao2023pytorch] is the first PyTorch-native ZeRO, following the same sharding format and limitation, but it is optimized in performance.
Table 1: Interleaved copy time (ms) compared to their corresponding collectives in FSDP2 for GPT-OSS-120B on 64 GPUs. Shard(0) is the default parameter sharding mode and Shard(1) is used when Shard(0) incurs large padding.
Gather side Reduce side
@@
fully_shard (FSDP2)   is the second PyTorch-native ZeRO, representing the state-of-the-art FSDP in the community. It replaces the concatenated shard design with per-parameter sharding, representing each tensor as a Shard(0) DTensor. This exposes maximal DTensor flexibility for FSDP parameters in communication, computation, and model checkpointing. However, this even sharding format is still far from enabling structure-aware training. Moreover, this per-parameter design introduces performance overhead from copying parameters in interleaved memory addresses, as shown in Figure 2 and Table 1.
Figure 2: The fundamental overhead in FSDP2. (AllGather is shown; ReduceScatter is a reverse process.)
Megatron-FSDP   [megatron_fsdp] is the most recent FSDP prototype that pursues speed. It forgoes FSDP2’s design and rolls back to FSDP1’s concatenated sharding to avoid the copying overhead, while heavily optimizing performance. However, Megatron-FSDP develops a special mechanism to enforce a concatenation-sharded tensor become a Shard(0) DTensor, such that the model checkpointing can use DTensor. This mechanism inserts padding into the concatenation so that tensors are sharded row-wise along device boundaries rather than element-wise. Without careful padding planning, the concatenation size can grow significantly, increasing both memory usage and communication volume. Moreover, row-wise sharding still falls short of supporting structure-aware training.
@@ 3 Overview
To address both challenges of flexibility and performance, we present veScale-FSDP, a novel FSDP that combines the best of worlds. Figure 3 provides the overview. Model developers are given the freedom to develop sophisticated large models (e.g., with sparse MoE structures) and structure-aware optimizers (e.g., with non-element-wise operators) for achieving unprecedented model quality. Meanwhile, the model/optimizer can be simply parallelized with PyTorch-native API fully_shard like FSDP2, without intrusively hacking model/optimizer code. During parallelization, complex operators of models/optimizers can still enjoy single-device semantics, thanks to a proposed sharding format, dubbed RaggedShard that offers the flexibility to express arbitrary sharding granularity and arbitrary distribution across devices for each DTensor (§4). Under the hood, RaggedShard DTensors are grouped for bucketed communication. Toward optimal performance, their layouts are rearranged via a proposed planning algorithm which is derived from a NP-hard optimization problem. The planned layouts are then mapped to a Distributed Buffer (DBuffer), a new primitive that achieves zero-copy and minimal overhead (§5), enabling efficient scaling up to 10K GPUs in real production deployments.
@@ 4 RaggedShard for Flexibility
This section proposes a novel and general sharding format RaggedShard to enable flexibility of FSDP for complex model and structure-aware optimizers.
@@
Existing sharding formats.   Figure 4 summarizes existing sharding formats for FSDP tensors. The most common yet conventional format is the Element-wise Shard, where tensors are partitioned arbitrarily across devices without respecting structural boundaries. This leads to dangling elements on each device and a loss of tensor shape and stride information. Consequently, this format severely limits flexibility: (i) non-element-wise operations such as matrix multiplications break on misaligned elements, (ii) collective communication for redistributing tensors across dimensions becomes complex and inefficient, and (iii) quantization methods such as block-wise FP8 quantization cannot align block sizes with random shard boundaries, introducing substantial communication overhead for cross-boundary synchronization. Unfortunately, this element-wise sharding design remains the backbone of DeepSpeed and FSDP1.
@@
The RaggedShard format.   Inspired by the JaggedTensor/NestedTensor on single device [jaggedtensor, nestedtensor, raggedtensors], we propose RaggedShard format for DTensor to offer the flexibility to express arbitrary sharding granularity in contiguous memory (the atomic non-shardable block consisting of contiguous elements or rows or planes) and arbitrary sharding distribution (the number of blocks per device). For a simple example in Figure 4, setting RaggedShard’s granularity as one tensor row gives Row-wise RaggedShard with different number of rows across devices. A similar concept has been prototyped in the model checkpointing mechanism of Megatron-FSDP.
@@
To support combinations of multiple parallelization strategies, RaggedShard needs to compose cleanly with existing DTensor placements. RaggedShard is orthogonal to both replicated and partial-value placements and veScale-FSDP specially handles the Shard placement. In practice, TP uses Shard(0) and Shard(1) for column- and row-wise tensor parallelism; EP can be encoded as Shard(0) along the expert dimension. By convention, EP/TP is applied before FSDP. In PyTorch, however, the DTensor placement list is organized in the opposite order of conceptual application (see Figure 5): a tensor shown with placements (RaggedShard, Shard(0)) is partitioned as Shard(0) followed by RaggedShard. veScale-FSDP reconciles this by: (i) for Shard(0), introducing a dedicated placement StridedRaggedShard that carries reordering/stride metadata and performs the required reshuffle when materializing the full tensor; and (ii) for Shard(dim>0), adapting the ragged sharding granularity so it never cuts into that dimension, choosing the granularity as Least Common Multiple (LCM): LCM(stride(dim), user-defined granularity).
Meanwhile, RaggedShard, as an extended DTensor placement, can offers the checkpointing capability by directly reusing DTensor-based checkpointing stacks (e.g., PyTorch Distributed Checkpoint [dcp]) for failure recovery and also inheriting their optimizations such as communication-free sharded checkpointing.
@@ 5 Grouped RaggedShard for Performance
This section discusses how to group RaggedShard DTensors for efficient communication, as well as its NP-hardness for optimality and the underlying optimization.
Figure 6: Grouped communication of RaggedShard DTensors.
Challenges for efficient communication.   As well known in the systems community, collective communication relies on tensor bucketing or grouping to maximize network utilization [li2020pytorch, zhao2023pytorch]. The same applies to RaggedShard DTensors in FSDP. However, efficiently grouping RaggedShard tensors is non-trivial and naively approaches can lead to significant inefficiencies. Figure 6(a) illustrates three inefficient factors: i) Sharded block that can happen when tensors are just concatenated back-to-back and put into communication buffer, without realizing that sharding boundary is within a certain block, which breaks the abstraction of Block-wise RaggedShard and incurs extra communication for quantization; ii) Non-contiguous tensor memory that can happen when two ends of communication buffers are padded (to align collective preferred unit size [wu2025terabyte, nccl16byte] or equal size across devices [nccl2025collective]), without realizing that such padding is within a certain tensor, which breaks the memory contiguousness and incurs interleaved copy overhead (e.g., similar to the copy-out after AllGather in Figure 2); and iii) Imbalanced load that can happen when different tensor sizes or block sizes or padding sizes are not aggregated to be equal across devices, which breaks the symmetry in collective communication (esp., Ring Algorithm) and ends up in underutilized networking.
@@
Optimization problem formulation.   Formally, the proposed approach can be formulated as an optimization problem. Let $\mathcal{T}={T_{1},T_{2},\ldots,T_{n}}$ denote a set of RaggedShard DTensors, which are sharded across $m$ devices. Each DTensor $t\in\mathcal{T}$ has a block size of $g_{t}$, a total tensor size (in elements) $e_{t}$, and hence $u_{t}=e_{t}/g_{t}$ sharding blocks. We allocate a global communication buffer and place each $t$ as a contiguous memory interval $\ell_{t},r_{t})$. The decision variables are the per-device buffer size $S$ and the interval endpoints $\{\ell_{t},r_{t}\}_{t\in\mathcal{T}}$. Our goal is to minimize $S$ subject to the three factor constraints: Non-Sharded Block, Contiguous Tensor Memory, and strict Balanced Load (Figure 6(b)):
@@
2:Output: minimal per-device buffer size $S$ .
3:
4:function CheckValidShard($S$)
5:  for all $t\in\mathcal{T}$ do
6:   $l\leftarrow 0$
7:   while $l<e_{t}$ do
8:    // Monotonic in i, skip intermediate calculation
9:    $r\leftarrow\max\{i\!\in\![l,e_{t}):dp(t,i;S)=dp(t,l;S)\}$
10:    Add interval $[l,r]$ into dp result.
11:    $l\leftarrow r+1$
12:   end while
13:  end for
14:  return: $dp(t_{last},e_{t_{last}};S)\leq m$
15:end function
@@
17:for all $g^{\prime}\in G$.sorted() do
18:  $g\leftarrow\text{LeastCommonMultiple}(g,g^{\prime})$
19:  $S^{\prime}\leftarrow\min\{k*g:k\in{N}\land$ CheckValidShard(kg)(k*g)}
20:  $S\leftarrow\min(S,S^{\prime})$
21:end for
22:return $S$
@@
Heuristic-guided solution.   veScale-FSDP proposes a polynomial-time dynamic-programming (DP) buffer-layout algorithm, guided by permutation heuristics that exploit the regularity of transformer models. The optimization difficulty comes from tensor permutation: in principle, any permutation of tensors could be mapped into the buffer, and finding the global optimum would require exploring all permutations. Fortunately, in practice, transformer parameters are highly structured: linear weights dominate the total parameter count, and sharding blocks are often consistent across layers. To leverage this regularity, we explore three simple permutations: (i) default order of tensors; (ii) sorted order by sharding block sizes, and (iii) sorted order by tensor shapes. Our statistics show that these orders yield optimal or near-optimal results, so we adopt the default order for simplicity and ease of debugging.
@@
Given the tensor order, the proposed algorithm applies a DP procedure to place tensors into a smallest global buffer while enforcing the aforementioned three factor constraints. It enjoys a time complexity of $O(|\mathcal{T}|^{2}m\log(E)\log(|\mathcal{T}|m))$. Algorithm 1 presents the detail.

### v3

# veScale-FSDP: Flexible and High-Performance FSDP at Scale

@@ 1 Introduction
Large language models (LLMs) have become a transformative technology in everyday applications. Driven by the scaling law [kaplan2020scaling], LLMs now reach billions of parameters and approach human-level performance in various domains. Training such giant models requires parallelization techniques that distribute the model and optimizer states across thousands of GPUs [jiang2024megascale]. Among these, Fully Sharded Data Parallel (FSDP) [zhao2023pytorch, pytorch2024fsdp2, megatron_fsdp], also known as DeepSpeed ZeRO [rajbhandari2020zero], is one of the most fundamental techniques. FSDP is often the first choice because of its memory efficiency and flexible data-parallel programming paradigm that is decoupled from model architecture. When additional scaling is needed, FSDP can be combined with other forms of parallelism [smith2022using, ma2025veomni].
@@
However, existing FSDP systems struggle to support modern structure-aware training methods. State-of-the-art models use non-element-wise optimizers such as Shampoo [gupta2018shampoo] and Muon [jordan2024muon], and block-wise quantized training as used in DeepSeek-V3 [liu2024deepseek], all of which require tensor blocks to remain intact under sharding. The core limitation is that existing FSDP frameworks shard model states either element-wise [rajbhandari2020zero, zhao2023pytorch] or row-wise [pytorch2024fsdp2, megatron_fsdp], producing sharding boundaries that often misalign with the required block structure. Consequently, either model developers must intrusively modify the model or optimizer code to match tensor boundaries, or system developers must handle complex boundary checks, padding, and additional communication logic.
@@
Beyond inflexibility, existing FSDP systems fall short of production throughput and memory targets, where we aim to extract every bit of hardware efficiency. GPU Memory is the tighter constraint: in shared clusters, jobs run out of memory or will operate at the memory limit incurring expensive device-side frees, prompting over-provisioning that leaves GPU resources wasted. These demands become even more critical when scaling training to over 10K GPUs and trillions of parameters. Few existing FSDP systems can scale to this level while maintaining efficiency. DeepSpeed ZeRO [rajbhandari2020zero] pioneered FSDP research but suffers from fragmented AllGather operations [deepspeed_AG] and inefficient memory management [fsdp_record_stream]. PyTorch FSDP1 [zhao2023pytorch] addresses some AllGather inefficiency, but incurs slow ReduceScatter [fsdp1_reduce] and does not solve memory overhead [fsdp_record_stream]. PyTorch FSDP2 [pytorch2024fsdp2] improves memory management [per_parameter_shard_rfc] but introduces a high tensor copy overhead. Meanwhile, both FSDP1 and FSDP2 suffer from slow collectives due to unaligned communication buffers [wu2025terabyte, nccl16byte]. Megatron-FSDP [megatron_fsdp] adopts a zero-copy design that avoids FSDP2’s tensor copy overhead but requires extra padding, increasing both communication and memory costs.
To address these flexibility and performance shortcomings, we present veScale-FSDP, a novel FSDP system that preserves the PyTorch-native fully_shard API while rearchitecting the PyTorch FSDP2 backend for flexible and high-performance training at scale:
@@
$\triangleright$ For performance, veScale-FSDP introduces a planning algorithm that rearranges RaggedShard tensors to maximize communication efficiency while respecting their desired sharding blocks. We formulate planning as an NP-hard optimization problem and use practical polynomial-time heuristics that achieve high-quality solutions in practice.
$\triangleright$ veScale-FSDP further provides a high-performance primitive, Distributed Buffer (DBuffer), that backs RaggedShard tensors with slices of a global buffer, enabling zero-copy access, reducing communication overhead, and reducing memory fragmentation via batched memory allocations.
@@
RaggedShard code is open-sourced at https://github.com/volcengine/veScale.
@@ 2.1 Structure–Aware Training
Structure-aware training underpins frontier models such as Kimi K2 [team2025kimi], Gemini [team2024gemini], and DeepSeek-V3 [liu2024deepseek], and is becoming increasingly important. Representative examples include:
Matrix Optimizers.   Matrix-based optimizers such as Shampoo [gupta2018shampoo] and Muon [jordan2024muon] can converge faster than AdamW [wen2025fantastic]. Their updates operate on parameters in their original 2D matrix rather than on independently sharded elements. In practice, this requires gathering each matrix on a selected device before performing the matrix-level update, and then redistributing the update back.
Block-wise Quantization.   Quantizing model weights [liu2024deepseek] and optimizer states [dettmers8] is widely used to improve training efficiency. Block-wise quantization is a common technique because per-block scaling factors mitigate the accuracy loss from reduced precision while preserving system efficiency. However, communication-free block-wise quantization requires each quantization block to reside entirely on a single device. If parameters are sharded without respecting block boundaries, devices must exchange scaling-factor metadata, complicating the implementation and reducing the efficiency gains.
@@ 2.2 DTensor and JaggedTensor
Figure 1: Distributed Tensor (DTensor) for flexible communication and computation. The figure shows an example of DTensors executing a sharded matrix multiplication (matmul) on a device. A row-sharded (Shard(0)) DTensor is first redistributed to a replicated (Replicate) DTensor so that it can be multiplied with a column-sharded (Shard(1)) DTensor, producing a column-sharded (Shard(1)) DTensor. The blue part in each DTensor indicates the materialized local tensor on that device.
Distributed Tensor (DTensor)   [xu2021gspmd, torch-dtensor, li2025vescale] is a promising primitive of PyTorch that provides an opportunity towards structure–aware training. It represents a logical global tensor distributed across multiple devices, with each device holding its local tensor. DTensor supports three sharding formats (placements): Shard(dim), which evenly shards a global tensor along a tensor dimension; Replicate, which replicates the global tensor on every device; and Partial, in which each device holds a partial value of the global tensor that must be reduced across devices to materialize the full result. It also provides a redistribute API that converts a DTensor between placements with implicit collective communications. Additionally, DTensors can be computed directly from operators such as matmul, as shown in Figure 1. However, a fundamental limitation remains for structure-aware training: the Shard format cannot represent the block-wise sharding needed for quantization or the uneven sharding required by matrix optimizers.
JaggedTensor/NestedTensor on a single device   [jaggedtensor, nestedtensor, raggedtensors] is a PyTorch/TensorFlow primitive for representing tensors whose last dimension is jagged. For example, a 2D tensor whose rows may have different lengths. These primitives cannot express sharding with block-level granularity, but they offer a useful hint for how veScale-FSDP can support structure-aware sharding in distributed training.
@@ 2.3 ZeRO and FSDPs
DeepSpeed ZeRO   [rajbhandari2020zero] pioneered this line of FSDP research. Its core idea is to concatenate a layer of tensors (parameters, gradients, and optimizer states) and then shard each concatenated tensor across devices, where tensors may be irregularly partitioned across device boundaries. ZeRO only unshards a layer using AllGather before the forward and backward passes, and reduces the layer gradients using ReduceScatter back across devices. Such a sharding design is fundamentally element-wise and cannot support structure-aware training.
FullyShardedDataParallel (FSDP1)   [zhao2023pytorch] is the first PyTorch-native ZeRO, following the same sharding format and limitation, but it is optimized for performance.
Table 1: Interleaved copy overhead in FSDP2 for GPT-OSS-120B on 64 H800 GPUs. The table reports the time of interleaved Copy-Out relative to AllGather on the AllGather path, and the time of interleaved Copy-In relative to ReduceScatter on the ReduceScatter path. Shard(0) is the default sharding format for each parameter, while Shard(1) is used when Shard(0) incurs large padding.
AllGather path ReduceScatter path
@@
fully_shard (FSDP2)   is the second PyTorch-native ZeRO and represents the state of the art in the FSDP community. It replaces the concatenated-shard design with per-parameter sharding, representing each tensor as a Shard(0) DTensor. This exposes DTensor flexibility for FSDP parameters in communication, computation, and checkpointing. However, this evenly sharded format still falls short of enabling structure-aware training. Moreover, FSDP2 introduces performance overheads by copying parameters into and from interleaved memory addresses, as shown in Figure 2 and Table 1.
Figure 2: The copy overhead in the FSDP2 design. Each parameter is represented as a Shard(0) DTensor and is evenly sharded across devices. Before AllGather, the input communication buffer on each device needs to copy in the local shards of all parameters. After AllGather, the output communication buffer stores each parameter in interleaved memory addresses rather than a contiguous address. As a result, FSDP2 must copy out each parameter from the output buffer to materialize the full tensor with a contiguous address needed for computation. The ReduceScatter path is a reversed one and incurs interleaved Copy-In overhead.
Megatron-FSDP   [megatron_fsdp] is the most recent FSDP prototype optimized for speed. It forgoes FSDP2’s design and reverts to FSDP1’s concatenated sharding to avoid copying overhead, while further optimizing performance. However, Megatron-FSDP introduces a special mechanism to make a concatenation-sharded tensor appear as a Shard(0) DTensor so that model checkpointing can reuse DTensor. This mechanism inserts padding into the concatenation so that tensors are sharded row-wise along device boundaries rather than element-wise. Without careful padding planning, the concatenation size can grow significantly, increasing both memory usage and communication volume. Moreover, row-wise sharding still falls short of supporting structure-aware training. Its granularity is fixed by the parameter layout and may not match what the block-wise quantization requires.
@@ 3 Overview
To address both flexibility and performance challenges, we present veScale-FSDP, a novel FSDP system. Figure 3 provides an overview. veScale-FSDP allows model developers to build sophisticated large models (e.g., with sparse MoE structures) and structure-aware optimizers (e.g., with non-element-wise operators) that achieve better model quality. At the same time, developers can parallelize these models and optimizers through PyTorch’s native fully_shard API, as in FSDP2, without intrusive modifications to model or optimizer code. During parallelization, complex model and optimizer operators can still preserve single-device semantics through a new sharding format, RaggedShard, which supports arbitrary sharding granularities over contiguous storage and arbitrary distributions across devices for each DTensor (§4). Under the hood, RaggedShard DTensors are grouped for bucketed communication. To optimize performance, their layouts are rearranged by a planning algorithm derived from an NP-hard optimization problem. The planned layouts are then mapped to a Distributed Buffer (DBuffer), a new primitive that enables zero-copy communication path with minimal overhead (§5). Together, these components enable veScale-FSDP to scale efficiently to 10K GPUs in real production deployments.
@@ 4 RaggedShard for Flexibility
This section proposes RaggedShard, a novel and general sharding format to increase FSDP’s flexibility for complex models and structure-aware optimizers.
@@
Existing sharding formats.   Figure 4 summarizes existing sharding formats for FSDP tensors. The most common and conventional format is the Element-wise Shard, where tensors are partitioned arbitrarily across devices without respecting structural boundaries. This leads to dangling elements on each device and a loss of tensor shape and stride information. Consequently, this format severely limits flexibility: (i) non-element-wise operations such as matrix multiplications break on misaligned elements, (ii) collective communication for redistributing tensors across dimensions becomes complex and inefficient, and (iii) quantization methods such as block-wise FP8 quantization cannot align block sizes with random shard boundaries, introducing substantial communication overhead for cross-boundary synchronization. Unfortunately, this element-wise sharding design remains the backbone of DeepSpeed and FSDP1.
@@
The RaggedShard format.   Inspired by JaggedTensor/NestedTensor [jaggedtensor, nestedtensor, raggedtensors], we propose the RaggedShard format for DTensor. RaggedShard increases flexibility by supporting arbitrary sharding granularity in contiguous memory, i.e., the size of the atomic non-shardable block, and arbitrary sharding distribution, i.e., the number of such blocks placed on each device. The sharding granularity can be defined over contiguous tensor units, such as elements, rows, or higher-dimensional planes. Figure 4 shows a simple example: when the sharding granularity is set to one tensor row, RaggedShard yields Row-wise RaggedShard, where different devices may hold different numbers of rows. A similar idea has been prototyped in the model checkpointing mechanism of Megatron-FSDP.
@@
To support combinations of multiple parallelization strategies, RaggedShard needs to compose cleanly with existing DTensor placements. RaggedShard is orthogonal to both replicated and partial-value placements and veScale-FSDP specially handles the Shard placement. In practice, TP uses Shard(0) and Shard(1) for column- and row-wise Tensor Parallelism; EP can be encoded as Shard(0) along the expert dimension. By convention, EP/TP is applied before FSDP. In PyTorch, however, the DTensor placement list is organized in the opposite order of conceptual application (see Figure 5): a tensor shown with placements (RaggedShard, Shard(0)) is partitioned as Shard(0) followed by RaggedShard. veScale-FSDP reconciles this by: (i) for Shard(0), introducing a dedicated placement StridedRaggedShard that carries reordering/stride metadata and performs the required reshuffle when materializing the full tensor; and (ii) for Shard(dim>0), adapting the ragged sharding granularity so it never cuts into that dimension by choosing the granularity as Least Common Multiple (LCM) of the tensor stride of that dimension and user-defined granularity.
Meanwhile, RaggedShard, as an extended DTensor placement, offers checkpointing capability by directly reusing DTensor-based checkpointing stacks (e.g., PyTorch Distributed Checkpoint [dcp]) for failure recovery and also inheriting their optimizations such as communication-free sharded checkpointing.
@@ 5 Grouped RaggedShard for Performance
This section discusses how to group RaggedShard DTensors for efficient communication. We formulate the underlying optimization problem, prove the NP-hardness of the problem, and present a polynomial-time heuristic algorithm that achieves high-quality solutions in practice.
Figure 6: Communication layouts for grouped RaggedShard DTensors. (a) shows a naive strategy that concatenates tensors directly into the communication buffer. This can split blocks across shard boundaries, introduce padding within tensors, and produce unequal per-device buffer sizes, leading to inefficient communication. (b) shows our structure-aware planner that carefully plans the communication layout – first permutes tensors and then inserts padding between tensors rather than within them. This preserves non-sharded blocks, keeps each tensor contiguous in memory, and balances per-device communication buffers, leading to efficient communication.
Challenges for efficient communication.   As is well known in the systems community, collective communication relies on tensor bucketing or grouping to maximize network utilization [li2020pytorch, zhao2023pytorch]. The same principle applies to communication over RaggedShard DTensors in veScale-FSDP. However, efficiently grouping RaggedShard tensors is non-trivial and straightforward approaches can lead to significant inefficiencies. Figure 6(a) illustrates three major sources of inefficiency:
  • Sharded block: Tensors are concatenated into a communication buffer without respecting block boundaries, causing individual blocks to be split across devices; this violates the abstraction of Block-wise RaggedShard and incurs additional communication for quantization.
  • Non-contiguous tensor memory: Padding may be inserted at the ends of communication buffers to satisfy collective alignment requirements [wu2025terabyte, nccl16byte] or equal-size constraints across devices [nccl2025collective], but may fall within a tensor. This breaks tensor contiguity and introduces interleaved copy overhead, similar to the Copy-Out after AllGather in Figure 2.
  • Imbalanced load: Differences in tensor sizes, block sizes, or padding sizes may lead to unequal buffer sizes across devices. This breaks the symmetry of collective communication and results in underutilized network bandwidth.
@@
Optimization problem formulation.   Formally, the proposed approach can be formulated as an optimization problem. Let $\mathcal{T}=\{t_{1},t_{2},\ldots,t_{n}\}$ denote a set of RaggedShard DTensors, which are sharded across $m$ devices. Each DTensor $t\in\mathcal{T}$ has a block size of $g_{t}$, a total tensor size (in elements) $e_{t}$, and therefore contains $u_{t}=e_{t}/g_{t}$ sharding blocks. We allocate a global communication buffer and place each $t$ in the buffer as a contiguous memory interval $\ell_{t},r_{t})$. The decision variables are a uniform per-device buffer size $S$ and the interval endpoints $\{\ell_{t},r_{t}\}_{t\in\mathcal{T}}$. The global communication buffer therefore has total size $mS$, partitioned into $m$ equal-sized device-local buffers, where device $k$ owns a memory interval of [(k1)S,kS)[(k-1)S,\,kS) for $k=1,\dots,m$. Our goal is to minimize $S$ subject to three constraints: Non-Sharded Block, Contiguous Tensor Memory, and Balanced Load (Figure 6(b)):
@@
2:Output: minimal uniform per-device buffer size $S^{\star}$.
3:Notation: for a candidate shard size $S$, $dp(t,i;S)$ denotes the minimum number of device-local shards required to place all tensors before $t$ and the first $i$ sharding blocks of tensor $t$.
4:
5:function CheckValidShard($S$)
6:  initialize segment records for $dp(\cdot,\cdot;S)$
7:  for all $t\in\mathcal{T}$ do
8:   $l\leftarrow 0$
9:   while $l<u_{t}$ do
10:    // $dp(t,i;S)$ is monotonic in $i$, so skip intermediate states within a constant segment
11:    $r\leftarrow\max\{i\!\in\![l,u_{t}):dp(t,i;S)=dp(t,l;S)\}$
12:    record segment $[l,r]$ with value $dp(t,l;S)$
13:    $l\leftarrow r+1$
14:   end while
15:  end for
16:  return $dp(t_{\mathrm{last}},u_{t_{\mathrm{last}}};S)\leq m$
17:end function
18:
@@
[…]
21:for all $g^{\prime}\in\textsc{SortAscending}(G)$ do
22:  $g\leftarrow\textsc{LeastCommonMultiple}(g,g^{\prime})$
23:  $S^{\prime}\leftarrow\min\{k*g:k\in{N}\land$ CheckValidShard(kg)(k*g)}
24:  $S^{\star}\leftarrow\min(S^{\star},S^{\prime})$
25:end for
26:return $S^{\star}$
@@
Heuristic-guided solution.   veScale-FSDP introduces a polynomial-time dynamic-programming (DP) buffer-layout algorithm, guided by permutation heuristics that exploit the regularity of transformer models. The optimization difficulty arises from tensor ordering: in principle, any permutation of tensors could be mapped into the global communication buffer, and finding the global optimum would require exploring all permutations. Fortunately, in practice, transformer parameters are highly structured: linear weights dominate the total parameter count, and sharding blocks are often consistent across layers. To leverage this regularity, we consider three tensor orders: (i) the default tensor order; (ii) sorting by sharding block size; and (iii) sorting by tensor shape. Our statistics show that these orders yield optimal or near-optimal results, so we adopt the default order for simplicity and ease of debugging. For non-transformer architectures, alternative tensor orderings can be plugged in without changing the DP algorithm itself.
@@
Given the tensor order, the proposed algorithm applies a DP procedure to place tensors into the smallest possible global buffer while enforcing the aforementioned three constraints. It has a time complexity of $O(|\mathcal{T}|^{2}m\log(E)\log(|\mathcal{T}|m))$. Algorithm 1 presents the details.